What are the mass % of C10H14O for the individual elements. Keep all answers to the first decimal place?


also, how many grams of oxygen would be present in a 375.3 g sample of this compound?


and how many oxygen atoms would that be?



1
Expert's answer
2021-03-09T06:15:58-0500

The molecular formula C10H14O has Molar Mass = 150.22 g/mol


Mass % = "\dfrac{mass of individual compound}{total mass of compound}"


Number of moles of compound = "\dfrac{375.3}{150.22}"


= 2.49 mol

Hence mass of oxygen = "2.49\times 16"


= 39.97 g


Number of oxygen atom = "39.97\times 6.022\times 10^{23}"


= "240.71\times 10^{23}"